OpenStudy (anonymous):

If y ≥ 1 then: a) y–2 ≥ 1 b) y–2 ≤ 1 c) 0 ≤ y–2 ≤ 1 d) 0 < y–2 ≤ 1 The answer is d but why is that? Also, what kind of question is this? Algebra? Inequalities?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

inequalities. not sure why its d though, sorry

OpenStudy (anonymous):

Thanks Ashley.

OpenStudy (anonymous):

@jim_thompson5910

sammixboo (sammixboo):

This question is an Inequality question, because it is dealing with inequalities, and idk why the answer is \(\color{lim}D\) , sorry! \(\bbox[5pt, olive ,border:5px solid #BAFF00]{\bf\huge \color{white}S\color{white}a\color{white}m\color{white}m\color{white}i\color{white}♥}\)

OpenStudy (broskishelleh):

@Zealous Which lesson and course is this for?

OpenStudy (imstuck):

That answer d really makes no sense at all. I'm with everyone else on that one!

OpenStudy (broskishelleh):

Ok @IMStuck You beat me to it this time!

OpenStudy (anonymous):

The answer should be: If (y is greater than or equal to 1), then ((y-2) is greater than or equal to -1).

OpenStudy (ikram002p):

y ≥ 1 y-2≥ -1 >0 sou can say that y-2 >0 but the part of y-2 \(\le 1\) is wrong

OpenStudy (ikram002p):

none of the choises is correct :) and i agree to @tangytaylor here a counter example that make D is incorrect let y=10 >1 y-2=10-2=8 >1 too its not less than 1